AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 289     Entry time: Sat Feb 26 23:20:13 2022
Author: David, Pierre-Michell 
Subject: Update status 

In this last hour we have been playing with the cryostat temperature as dictated by Nikos, in order to have the target working properly. For this, we have increased from 80K to 140K the cryostat temperature then we saw again target until we didn't. Then we went down again to 80K. After a few minutes, with a really nice target again was gone, so we increased the temperature to 120K at 00:19. At 00:27 decrease from 120K to 80K.

Ring current 8e7 particles/s

Merger, Tape server (WS3): 1580 kB/s writing ok

Leakage current (WS4): 3.27 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 2022-02-27_00-26-44xytotal.png  187 kB  | Hide | Hide all
2022-02-27_00-26-44xytotal.png
Attachment 2: 2022-02-27_00-26-22exrate.png  155 kB  | Hide | Hide all
2022-02-27_00-26-22exrate.png
Attachment 3: 2022-02-27_00-25-29extotal.png  163 kB  | Hide | Hide all
2022-02-27_00-25-29extotal.png
Attachment 4: 2022-02-27_00-25-07temp.png  111 kB  | Hide | Hide all
2022-02-27_00-25-07temp.png
Attachment 5: 2022-02-27_00-24-47goodevents.png  102 kB  | Hide | Hide all
2022-02-27_00-24-47goodevents.png
Attachment 6: 2022-02-27_00-23-42current.png  20 kB  | Hide | Hide all
2022-02-27_00-23-42current.png
Attachment 7: 2022-02-27_00-31-26.png  165 kB  | Hide | Hide all
2022-02-27_00-31-26.png
ELOG V3.1.4-unknown